# Randomized Space–Time Coded Stacked Intelligent Metasurfaces for Massive Multiuser Downlink Connectivity

### v1

@@ V Transmission Coefficient Synthesis for randomized ST-SIM
Step 2 - Generation of the matrix ${\mathbf{\Delta}}(t)$: The transmission coefficient $\delta_{z}(t)$ of the $z$-th meta-atom in the initial DAL, with $z\in\mathcal{Z}$, is generated according to (5). Specifically, at the beginning of each time-slot, we choose the digitally controllable phases $\psi^{(m)}_{z}$ of the transmission coefficients $\delta^{(m)}_{z}=\alpha_{\text{dal}}\,e^{j\psi^{(m)}_{z}}$ in (5), for $m\in\mathcal{M}$ and $z\in\mathcal{Z}$, as a sequence of i.i.d. random variables with respect to both $m$ and $z$, where each random variable $\psi^{(m)}_{z}$ is uniformly distributed in the interval $0,2\pi)$. By varying the phases of the meta-atoms in the first layer across both space and time, the SIM implements a ST-coded beamforming $\widetilde{{\mathbf{G}}}(t)$ that randomly generates $N$ steering vectors. At this point, relying only on partial CSIT, the transmitter schedules the transmission towards the $N$ users that are closest to the resulting beams, which are more likely to ensure signal power maximization as the total number of users $U$ increases. This fact is corroborated by the numerical results shown in the next section.
@@ VI Numerical results
In this section, we present Monte Carlo simulations to validate the proposed ST-SIM architecture and assess its achievable sum-rate capacity in multiuser downlink.
Figure 3: Objective function (42) versus the number of passive layers $L_{\text{pc}}$ for $Q\in\{25,36,49,64\}$ meta-atoms. Parameters are set to $Z=9$, $V=25$, and $L_{\text{ac}}=4$. All metasurface layers are square. The signal from the RF chain first passes through the AC layers and, then, through the PC layers.
To evaluate the effectiveness of incorporating DALs into the proposed SIM architecture, we consider a conventional space-only SIM in which the first DAL is S-coded, too, like the other (L1)(L-1) layers, i.e., $\boldsymbol{\Delta}(t)=\boldsymbol{\Gamma}_{1}\triangleq{\text{diag}}(\boldsymbol{\gamma}_{1})\in\mathbb{C}^{Z}$. Accordingly, the SIM response (15) becomes
${\mathbf{G}}={\mathbf{\Gamma}}_{\text{dal}}\,{\mathbf{W}}_{L}\,{\mathbf{\Gamma}}_{L-1}\,{\mathbf{W}}_{L-1}\cdots{\mathbf{\Gamma}}_{2}\,{\mathbf{W}}_{2}\,\boldsymbol{\Gamma}_{1}\in\mathbb{C}^{V\times Z}\>.$ (57)
In this case, we optimize the objective function (42), where ${\mathbf{G}}_{0}$ is replaced with (57) and the target matrix ${\mathbf{G}}_{\text{targ}}\in\mathbb{C}^{V\times Z}$, which represents the desired forward transfer function from the transmit UPA to the output of the SIM, is constructed as
${\mathbf{G}}_{\text{targ}}=\frac{1}{\beta\,\|{\mathbf{W}}_{1}\|}\,{\mathbf{R}}^{\text{H}}_{\text{a}}\left({\mathbf{R}}_{\text{a}}\,{\mathbf{R}}^{\text{H}}_{\text{a}}\right)^{-1/2}$ (58)
where ${\mathbf{R}}_{\text{a}}\in\mathbb{C}^{Z\times V}$ is a random matrix with entries independently drawn from a circularly symmetric complex Gaussian distribution with zero mean and unit variance. By construction, the columns of ${\mathbf{G}}_{\text{targ}}$ are orthogonal and fulfill (25).
In this setup, the two boundary DALs use $Z_{x}=Z_{y}=3$ and $V_{x}=V_{y}=5$ meta-atoms, respectively, and the number of AC layers is $L_{\text{ac}}=4$. The AC layers are placed at the input of the SIM stack, so the signal from the RF chain first passes through the AC layers and then through the PC layers. The placement of AC/PC layers, indeed, affects the convergence rate of the PDG algorithm, which achieves its fastest convergence when the AC layers are located before the PC ones [18].
Figure 4: Convergence rate of the PGD algorithm for $Q\in\{25,36,49,64\}$ meta-atoms. Parameters are set to $Z=9$, $V=25$, $L_{\text{ac}}=4$, and $L_{\text{pc}}=8$. All metasurface layers are square. The signal from the RF chain first passes through the AC layers and, then, through the PC layers.
@@
Figure 3 shows the value of the objective function (42) plotted as a function of the number of PC layers $L_{\text{pc}}$, ranging from $L_{\text{pc}}=4$ to $L_{\text{pc}}=14$, for different values of the number of meta-atoms $Q$. The PC layers are assumed to have constant transmittance $\alpha_{\text{pc}}=0.9$. The case $Q=V=25$ meta-atoms corresponds to the baseline SIM architecture of size $Z\times Q$, where the final DAL is not employed and, thus, the only design variable available for synthesizing the target matrix ${\mathbf{G}}_{\text{targ}}\in\mathbb{C}^{Q\times Z}$ is the number of passive layers $L_{\text{pc}}$. It can be observed from Fig. 3 that, for a fixed number of meta-atoms $Q$, the performance improves as the number of layers increases. More interestingly, however, one can infer from the same figure that, for a fixed number of passive layers $L_{\text{pc}}$, the objective function (42) also improves with increasing $Q$, provided that DAL is employed. This additional design flexibility, indeed, enables the system to achieve arbitrarily small squared errors $\|{\mathbf{G}}_{0}-{\mathbf{G}}_{\text{targ}}\|^{2}$. On the contrary, beyond a certain number of layers, the objective function (42) no longer decreases monotonically for a conventional SIM, due to error propagation effects across layers in the PGD algorithm [9].
@@
Figure 4 presents the convergence rate of the PGD algorithm for $Q\in\{25,36,49,64\}$, with a fixed number of passive layers $L_{\text{pc}}=8$. As expected, the objective function $\|{\mathbf{G}}_{0}-{\mathbf{G}}_{\text{targ}}\|^{2}$ decreases with iterations for all configurations. More importantly, this figure shows that increasing $Q$ significantly improves the convergence performance. Specifically, higher values of $Q$ yield both faster convergence and lower steady-state error, allowing the objective function to reach values approximately below $25\,\text{dB}$. This fact confirms that increasing the number of meta-atoms, and thereby increasing the available degrees of freedom in the DAL-aided SIM synthesis, enables more accurate approximation of the target matrix ${\mathbf{G}}_{\text{targ}}$. The case $Q=25$ meta-atoms corresponds to the baseline SIM configuration (i.e., without DAL), which achieves the worst performance among the tested architectures. In contrast, larger values of $Q$ provide a substantial gain, thanks to the additional absorbing meta-atoms introduced by the DAL.
We assess now the sum-rate capacity achieved by the proposed randomized ST-SIM-based multiuser downlink. We consider a three-dimensional Cartesian coordinate system wherein the base station is positioned at $(0,0,h_{\text{BS}})$, with $h_{\text{BS}}=10$ m. The locations of the $U$ users are modeled as random variables uniformly distributed within a circular annulus of inner radius $r_{\text{i}}=10$ m and outer radius $r_{\text{o}}=50$ m, centered at $(0,0,0)$ in the $xy$-plane. The system operates at carrier frequency $f_{0}=28$ GHz, with a transmission bandwidth of $10$ MHz and a noise power spectral density of 174-174 dBm/Hz for all users. The total available power budget at the BS, including transmit array gain, is fixed at $\mathscr{P}_{\text{rad}}=15$ dBm. The BS is equipped with a uniform planar array with inter-element spacing $d_{\text{meta}}={\lambda_{0}}/{2}$ and $N_{x}=N_{y}=2$ antennas (i.e., $N=4$). The ST-SIM layers are uniformly spaced by a distance of $s_{\text{lay}}=\lambda_{0}/2$. The ST-coded initial DAL uses $Z_{x}=Z_{y}=10$ meta-atoms, while the S-coded terminal DAL $\boldsymbol{\Gamma}_{\text{dal}}$ employs $V_{x}=V_{y}=3$. The intermediate layers have $Q_{x}=Q_{y}=24$ meta-atoms, and the numbers of PC and AC layers are set to $L_{\text{pc}}=6$ and $L_{\text{ac}}=2$, respectively. The number of time-slots is $M=2$.
Figure 5 reports the TA sum-rate capacity (36) as a function of the number of users $U$ for the proposed randomized ST-SIM. For comparison, we also consider a conventional MIMO scheme in which each user feeds back to the transmitter its channel vector $\mathbf{h}_{u}$. With full CSI at the transmitter, the BS can implement optimal linear beamforming (e.g., sum-rate-maximizing or zero-forcing), but selecting (at the transmitter) the optimal subset of users from the $U$ candidates. However, such an approach entails a combinatorial search with a computational burden that is practical only for relatively small user pools [18]. When the number of users $U$ is large, a practical approach [29] is to align transmission with the channels of the $N$ users exhibiting the highest $\|{\mathbf{h}}_{u}\|^{2}$, with $u\in\mathcal{U}$. Specifically, the $N$ data streams are linearly precoded using the steering vectors $\widetilde{{\mathbf{g}}}_{{u}^{\star}_{n}}=\tfrac{h_{{u}^{\star}_{n}}}{\|{\mathbf{h}}_{{u}^{\star}_{n}}\|^{2}}$, with $u^{\star}_{n}\in\mathcal{U}^{\star}$}, and then transmitted by the $V$ antennas, where $\mathcal{U}^{\star}$ is the index set of the selected users.
@@
Figure 5: TA sum-rate capacity as a function of the number of users $U$. The ST-SIM parameters are $Q=576$, $Z=100$, $V=9$, $N=4$, $L_{\text{ac}}=2$, and $L_{\text{pc}}=6$. All metasurface layers are square. The signal from the RF chain first passes through the AC layers and, then, through the PC layers.
It is evident from Fig. 5 that, as the number of users $U$ increases, the sum-rate capacity of the proposed scheme grows and, for $U>200$, exceeds that of the conventional MIMO scheme, which does not exploit multiuser diversity. It is worth noting that, for $U\leq 200$ users, the superiority of the conventional MIMO scheme stems from the fact that the transmitter has perfect knowledge of the individual channel amplitude and phase for all antennas, which requires a substantially higher feedback than of the proposed randomized ST-SIM (see Remark 3).

### v2

@@ V Transmission Coefficient Synthesis for randomized ST-SIM
Step 2 - Generation of the matrix ${\mathbf{\Delta}}(t)$: The transmission coefficient $\delta_{z}(t)$ of the $z$-th meta-atom in the initial DAL, with $z\in\mathcal{Z}$, is generated according to (5). Specifically, at the beginning of each time-slot, we choose the digitally controllable phases $\psi^{(m)}_{z}$ of the transmission coefficients $\delta^{(m)}_{z}=\alpha_{\text{dal}}\,e^{j\psi^{(m)}_{z}}$ in (5), for $m\in\mathcal{M}$ and $z\in\mathcal{Z}$, as a sequence of i.i.d. random variables with respect to both $m$ and $z$, where each random variable $\psi^{(m)}_{z}$ is uniformly distributed in the interval $0,2\pi)$. By varying the phases of the meta-atoms in the first layer across both space and time, the SIM implements a ST beamforming $\widetilde{{\mathbf{G}}}(t)$ that randomly generates $N$ steering vectors. At this point, relying only on partial CSIT, the transmitter schedules the transmission towards the $N$ users that are closest to the resulting beams, which are more likely to ensure signal power maximization as the total number of users $U$ increases. This fact is corroborated by the numerical results shown in the next section.
Algorithm 1 summarizes the pseudo-code of the DAL-aided ST-SIM synthesis procedure. It is important to distinguish between the offline block-level synthesis cost and the online slot-level update cost. The former is dominated by the PGD design of the S-only block and is incurred once per coherence interval, whereas the latter only involves random phase refresh of the ST layer and simple user scheduling operations.
Remark 4: The proposed ST-SIM-based randomization should not be confused with classical random beamforming techniques, such as those introduced in opportunistic beamforming frameworks [42, 38]. In these schemes, randomness is introduced at the digital baseband level by generating random precoding or beamforming vectors, while the physical wireless channel remains invariant over each coherence interval. In contrast, the proposed ST-SIM framework performs randomization directly in the wave domain through a ST metasurface layer that physically modulates the EM wavefront. This mechanism induces artificial time variations in the effective channel itself, even under slowly varying propagation conditions, thereby enabling multiuser diversity through physically induced channel fluctuations rather than digital precoder randomness. As a result, opportunistic scheduling can be realized with partial CSIT and reduced reliance on high-dimensional digital beamforming architectures.
Remark 5: The proposed ST-SIM-based randomization is fundamentally different from conventional multi-beam training or beam-sweeping techniques used for beam alignment and channel acquisition [42, 43]. Beam-sweeping methods sequentially transmit a predefined set of deterministic beams from a finite codebook with the primary aim of identifying favorable propagation directions and estimating channel parameters. In contrast, the proposed ST-SIM framework is not intended for beam discovery or alignment. Instead, it introduces randomized ST wave-domain transformations via a time-modulated metasurface layer, which physically induce artificial time variations in the effective channel. This mechanism enables opportunistic scheduling and multiuser diversity under partial CSIT, without relying on explicit beam indexing, deterministic beam codebooks, or beam selection procedures.
Remark 6: Besides the adiabatic condition $f_{s}\ll f_{0}$, a practically relevant requirement is that the settling time of the ST layer be sufficiently shorter than the slot duration $T_{s}$, so that the metasurface reaches its new configuration before payload transmission in the corresponding slot. For instance, with $T=5$ ms and $M=4$, one has $T_{s}=1.25$ ms and hence $f_{s}=800$ Hz, i.e., a reconfiguration speed on the order of kHz. Therefore, the proposed scheme does not require symbol-rate switching, since the ST coefficients are updated only at slot boundaries. This requirement is compatible with fast electronically tunable implementations, such as PIN-diode- or varactor-based metasurfaces [28], whereas slower technologies such as liquid-crystal-based metasurfaces [44] may require smaller values of $M$ or longer channel coherence times.
Algorithm 1 Synthesis of the DAL-aided ST-SIM
1:Target matrix ${\mathbf{G}}_{\mathrm{targ}}\in\mathbb{C}^{V\times Z}$; propagation matrices $\{{\mathbf{W}}_{\ell}\}$; index sets $\mathcal{L}^{(s)}=\mathcal{L}^{(s)}_{\mathrm{pc}}\cup\mathcal{L}^{(s)}_{\mathrm{ac}}$; amplitude bounds $\alpha_{\min},\alpha_{\max}]$; maximum number of iterations $\kappa_{\max}$; number of time-slots $M$.
2:Time-invariant response ${\mathbf{G}}_{0}$ and time-varying randomization matrices $\{{\mathbf{\Delta}}[m]\}_{m=0}^{M-1}$.
3:Step 1: Synthesis of ${\mathbf{G}}_{0}$ (performed once at the beginning of the algorithm)
4:Initialize transmission coefficients $\boldsymbol{\gamma}_{\ell}^{(0)}$ for $\ell\in\mathcal{L}^{(s)}$
5:for $\kappa=0,1,\ldots,\kappa_{\max}-1$ do
6:  for $\ell\in\mathcal{L}^{(s)}$ do
7:   ${\mathbf{E}}_{\ell}\triangleq{\mathbf{\Gamma}}_{\mathrm{dal}}{\mathbf{W}}_{L}{\mathbf{\Gamma}}_{L-1}{\mathbf{W}}_{L-1}\cdots{\mathbf{\Gamma}}_{\ell+1}{\mathbf{W}}_{\ell+1}$
8:   ${\mathbf{B}}_{\ell}\triangleq{\mathbf{W}}_{\ell}{\mathbf{\Gamma}}_{\ell-1}{\mathbf{W}}_{\ell-1}\cdots{\mathbf{\Gamma}}_{2}{\mathbf{W}}_{2}$
9:   ${\mathbf{A}}_{\ell}\triangleq({\mathbf{B}}_{\ell}^{\ast}{\mathbf{B}}_{\ell}^{\top})\circ({\mathbf{E}}_{\ell}^{\mathrm{H}}{\mathbf{E}}_{\ell})$
10:   ${\mathbf{v}}_{\ell}\triangleq\Big({\mathbf{E}}_{\ell}^{\mathrm{H}}\circ({\mathbf{B}}_{\ell}^{\ast}{\mathbf{G}}_{\mathrm{targ}}^{\top})\Big){\mathbf{1}}_{Q}$
11:   Compute the difference ${\mathbf{r}}_{\ell}\triangleq{\mathbf{A}}_{\ell}\,\boldsymbol{\gamma}_{\ell}-{\mathbf{v}}_{\ell}$
12:   if $\ell\in\mathcal{L}^{(s)}_{\mathrm{pc}}$ then
13:     Compute $\nabla_{\boldsymbol{\phi}_{\ell}}f=2\,\Im\big\{\mathrm{diag}(\boldsymbol{\gamma}_{\ell}^{\ast})\,{\mathbf{r}}_{\ell}\big\}$
14:     Update $\boldsymbol{\phi}_{\ell}\leftarrow\boldsymbol{\phi}_{\ell}-\mu_{\ell}\nabla_{\boldsymbol{\phi}_{\ell}}f$
15:     Project onto the unit circle $\boldsymbol{\gamma}_{\ell}\leftarrow\exp\!\big(j\boldsymbol{\phi}_{\ell}\big)$
16:   else if $\ell\in\mathcal{L}^{(s)}_{\mathrm{ac}}$ then
17:     Compute $\nabla_{\boldsymbol{\alpha}_{\ell}}f=2\,\Re\big\{\mathrm{diag}(\boldsymbol{\gamma}_{\ell}^{\ast})\,{\mathbf{r}}_{\ell}\big\}$
18:     Update $\boldsymbol{\alpha}_{\ell}\leftarrow\mathcal{P}_{[\alpha_{\min},\alpha_{\max}]}\!\left(\boldsymbol{\alpha}_{\ell}-\mu_{\alpha,\ell}\nabla_{\boldsymbol{\alpha}_{\ell}}f\right)$
19:     Set $\boldsymbol{\gamma}_{\ell}\leftarrow\boldsymbol{\alpha}_{\ell}\odot\exp\!\big(j\boldsymbol{\phi}_{\ell}\big)$
20:   end if
21:   Update ${\mathbf{G}}_{0}\leftarrow{\mathbf{\Gamma}}_{\mathrm{dal}}{\mathbf{W}}_{L}{\mathbf{\Gamma}}_{L-1}{\mathbf{W}}_{L-1}\cdots{\mathbf{\Gamma}}_{2}{\mathbf{W}}_{2}$
22:  end for
23:end for
24:Step 2: ST randomization of the initial DAL (performed once per time-slot)
25:for $m=0,1,\ldots,M-1$ do
26:  Draw i.i.d. random phases $\psi_{z}[m]$
27:  Set: $\delta_{z}[m]\triangleq e^{j\psi_{z}[m]}$
28:  Set: $\boldsymbol{\Delta}[m]\triangleq\mathrm{diag}\big(\delta_{0}[m],\ldots,\delta_{Z-1}[m]\big)$
29:  Compute: ${\mathbf{G}}[m]\triangleq{\mathbf{G}}_{0}\,{\mathbf{\Delta}}[m]$
30:end for
@@ VI Numerical results
In this section, we present Monte Carlo simulations to validate the proposed ST-SIM architecture and evaluate its achievable sum-rate in a multiuser downlink scenario. We first analyze an S-only (time-invariant) SIM configuration to quantify the impact of DAL on the synthesis accuracy, convergence behavior, and power efficiency of the multilayer structure through the objective function in (44). We then consider the full randomized ST-SIM framework with partial CSIT and compare it with a full-CSIT beamforming benchmark, evaluating the time-averaged sum-rate and fairness as functions of the user population and the number of time slots.
Figure 3: Objective function (44) versus the number of passive layers $L_{\text{pc}}$ for $Q\in\{25,36,49,64\}$ meta-atoms. Parameters are set to $Z=9$, $V=25$, and $L_{\text{ac}}=4$. The case $Q=V=25$ corresponds to a baseline SIM architecture without absorbing meta-atoms.. All metasurface layers are square. The signal from the RF chain first passes through the AC layers and, then, through the PC layers.
VI-A DAL-aided vs baseline S-only SIM synthesis
To assess the effectiveness of incorporating the DAL, which introduces QVQ-V perfectly absorbing elements, we compare the proposed SIM architecture with a conventional configuration where $Q=V$ and, thus, all meta-atoms are transmitting. To this end, we consider an S-only SIM configuration whose response reduces to (16). In this setting, we optimize (44), where the target matrix ${\mathbf{G}}_{\text{targ}}\in\mathbb{C}^{V\times Z}$, representing the desired block response, is constructed as a Haar matrix
${\mathbf{G}}_{\text{targ}}=\frac{\sqrt{N}}{\beta\,\|{\mathbf{W}}_{1}\|}\,{\mathbf{R}}^{\text{H}}_{\text{a}}\left({\mathbf{R}}_{\text{a}}\,{\mathbf{R}}^{\text{H}}_{\text{a}}\right)^{-1/2}$ (59)
where ${\mathbf{R}}_{\text{a}}\in\mathbb{C}^{Z\times V}$ is a random matrix whose entries are independently drawn from a circularly symmetric complex Gaussian distribution with zero mean and unit variance. By construction, the columns of ${\mathbf{G}}_{\text{targ}}$ fulfill (25). This choice prevents favoring particular spatial directions and allows assessing the synthesis capability of the proposed architecture in a general setting. In this setup, we set $Z_{x}=Z_{y}=3$ (i.e., $Z=9$), $V_{x}=V_{y}=5$ (i.e., $V=25$), and $L_{\text{ac}}=4$. The AC layers are placed at the input of the SIM stack, so the signal from the RF chain first passes through the AC layers and then through the PC layers. The placement of AC/PC layers, indeed, affects the convergence rate of the PDG algorithm, which achieves its fastest convergence when the AC layers are located before the PC ones [18].
Figure 4: Convergence rate of the PGD algorithm for $Q\in\{25,36,49,64\}$ meta-atoms. Parameters are set to $Z=9$, $V=25$, $L_{\text{ac}}=4$, and $L_{\text{pc}}=8$. The case $Q=V=25$ corresponds to the baseline SIM without DAL. All metasurface layers are square. The signal from the RF chain first passes through the AC layers and, then, through the PC layers.
@@ VI-A DAL-aided vs baseline S-only SIM synthesis
Figure 3 reports the objective function (44) as a function of the number of PC layers $L_{\mathrm{pc}}$, ranging from 4 to 14, for different values of the number of meta-atoms $Q$. The PC layers are assumed to have constant transmittance $\alpha_{\mathrm{pc}}=0.9$. When $Q=V=25$, the considered transmit structure reduces to a baseline SIM architecture of size $Z\times Q$ with no absorbing meta-atoms. In this case, the synthesis of the target matrix ${\mathbf{G}}_{\text{targ}}\in\mathbb{C}^{Q\times Z}$ depends only on the number of passive layers $L_{\mathrm{pc}}$. By contrast, in the DAL-aided SIM both $L_{\mathrm{pc}}$ and $Q$ act as design degrees of freedom. As shown in Fig. 3, the DAL-aided SIM with QVQ-V absorbing meta-atoms consistently outperforms the baseline configuration for all values of $L_{\mathrm{pc}}$. For a fixed number of meta-atoms $Q$, the performance improves as the number of PC layers increases. Moreover, for a fixed number of passive layers $L_{\mathrm{pc}}$, the objective function (44) decreases as $Q$ increases when the DAL is employed ($Q>V$). This additional flexibility enables arbitrarily small approximation errors $\|{\mathbf{G}}_{0}-{\mathbf{G}}_{\text{targ}}\|^{2}$. Conversely, in the baseline SIM the objective function ceases to decrease monotonically beyond a certain number of layers due to error propagation in the PGD iterations [9].
@@
Figure 4 shows the convergence behavior of the PGD algorithm for $Q\in\{25,36,49,64\}$ with $L_{\mathrm{pc}}=8$. The objective function $\|{\mathbf{G}}_{0}-{\mathbf{G}}_{\text{targ}}\|^{2}$ decreases with the number of iterations for all configurations. Increasing $Q$ significantly improves convergence, yielding both faster decay and lower steady-state error, with values below approximately $25\,\mathrm{dB}$. This confirms that enlarging the number of meta-atoms increases the available degrees of freedom in the DAL-aided synthesis and improves the approximation of ${\mathbf{G}}_{\text{targ}}$. The case $Q=25$ corresponds to the baseline SIM (without DAL) and exhibits the worst performance, whereas larger values of $Q$ provide substantial gains due to the additional absorbing meta-atoms introduced by the DAL.
Figure 5: Power gain of a DAL-aided and conventional SIM employing only PC layers for $V\in\{9,16,25\}$ meta-atoms. Parameters are set to $Z=9$ and $Q=25$. The case $Q=V=25$ corresponds to a baseline SIM architecture without absorbing meta-atoms. All metasurface layers are square.
We now analyze the impact of the DAL on the power efficiency of the SIM. The presence of absorbing meta-atoms introduces intrinsic power dissipation, since part of the incident EM energy is absorbed rather than transferred through the multilayer structure. Therefore, it is of interest to quantify how efficiently the SIM transfers power from input to output and compare it with a baseline configuration where all meta-atoms are transmitting. In this respect, we define the power gain of the SIM as
$G_{\mathrm{P}}\triangleq\frac{\mathscr{P}_{\mathrm{rad}}}{\mathscr{P}_{\mathrm{sig}}}$ (60)
where $\mathscr{P}_{\mathrm{rad}}$ is the radiated power in (24) and $\mathscr{P}_{\mathrm{sig}}$ is the input signal power. Although the total energy efficiency also depends on the consumption of the active layers, this contribution is implementation-dependent. Hence, we focus on the technology-agnostic metric $G_{\mathrm{P}}$. From (24), the radiated power becomes $\mathscr{P}_{\mathrm{rad}}=(\mathscr{P}_{\mathrm{sig}}/N)\|{\mathbf{G}}_{0}\|^{2}$, hence yielding
$G_{\mathrm{P}}=\frac{\|{\mathbf{G}}_{0}\|^{2}}{N}\>.$ (61)
A purely passive phase-only SIM generally exhibits a gain significantly below the bound implied by (25) due to propagation losses, whereas the inclusion of amplitude-controlled (active) layers allows approaching this limit through local amplification. The analytical evaluation of $G_{\mathrm{P}}$ is complicated by the cascaded products of diagonal transmission matrices and coupling matrices. Therefore, we rely on numerical results to compare the DAL-aided and baseline architectures.
Figure 6: Objective function (44) of a DAL-aided and conventional SIM employing only PC layers versus $L_{\text{pc}}$. Parameters are set to $Z=9$ and $Q=25$. The case $Q=V=25$ corresponds to a baseline SIM architecture without absorbing meta-atoms. All metasurface layers are square.
We consider $Z=9$, $Q=25$, and $V\in\{9,16,25\}$, where $V=Q$ corresponds to the baseline SIM. Figure 5 reports the power gain for a purely passive configuration employing only PC layers as a function of $L_{\mathrm{pc}}$. The baseline SIM achieves a higher power gain than the DAL-aided configurations, and the gap increases as the number of absorbing meta-atoms grows (i.e., decreasing $V$). However, this comparison does not account for synthesis accuracy. As shown in Fig. 6, the baseline SIM reaches a squared error comparable to the DAL-aided architecture only for $L_{\mathrm{pc}}\geq 16$, whereas the DAL-aided SIM attains low error with significantly fewer layers. Hence, in passive implementations the DAL trades a moderate reduction in power gain for a substantially lower number of layers required for convergence.
@@
Figure 7: Power gain of a DAL-aided and conventional SIM employing PC and AC layers for $V\in\{9,16,25\}$ meta-atoms. Parameters are set to $Z=9$, $Q=25$, and $L_{\text{ac}}=4$. The case $Q=V=25$ corresponds to a baseline SIM architecture without absorbing meta-atoms. All metasurface layers are square. The signal from the RF chain first passes through the AC layers and, then, through the PC layers.
Figure 7 shows the power gain when both passive and active layers are used ($L_{\mathrm{ac}}=4$). In this case, the DAL-aided architectures ($V<Q$) outperform the baseline SIM for all values of $L_{\mathrm{pc}}$, with the best performance obtained for $V=9$. This behavior indicates that, in the presence of active layers, the additional degrees of freedom introduced by the DAL improve not only the approximation accuracy but also the capability of concentrating power on the transmitting meta-atoms. This observation is consistent with the objective-function behavior shown in Fig. 8.
VI-B Multiuser downlink performance of the ST-SIM
Herein, we evaluate the effective sum-rate achieved by the proposed randomized ST-SIM multiuser downlink. The BS is located at $(0,0,h_{\mathrm{BS}})$ with $h_{\mathrm{BS}}=10$ m. The $U$ users are uniformly distributed in a circular annulus centered at the origin on the $xy$-plane, with inner radius $r_{i}=10$ m and outer radius $r_{o}=100$ m. The channel variance toward the $u$-th user is $\sigma_{u}^{2}=d_{u}^{-\eta}\lambda_{0}^{2}/(4\pi)^{2}$, where $d_{u}$ denotes the link distance and $\eta=1.6$. The coherence time is $T=5$ ms. The carrier frequency is $f_{0}=28$ GHz with bandwidth $10$ MHz, roll-off factor $\rho_{s}=0.25$, and noise spectral density 174-174 dBm/Hz. The total radiated power is $\mathscr{P}_{\mathrm{rad}}=15$ dBm. The BS employs a UPA with spacing $d_{\mathrm{meta}}=\lambda_{0}/2$ and $N_{x}=N_{y}=2$ ($N=4$). The ST-SIM layers are spaced by $s_{\mathrm{lay}}=\lambda_{0}/2$. The initial DAL uses $Z_{x}=Z_{y}=3$ meta-atoms, intermediate layers use $Q_{x}=Q_{y}=24$, and the number of PC and AC layers is $L_{\mathrm{pc}}=6$ and $L_{\mathrm{ac}}=2$, respectively.
As benchmark, we consider a fully-digital MIMO system with full CSIT, where each user feeds back its channel vector ${\mathbf{h}}_{u}$. The transmitter selects $N$ users and applies linear precoding using normalized channel directions $\widetilde{{\mathbf{g}}}_{u_{n}^{\star}}={\mathbf{h}}_{u_{n}^{\star}}/\|{\mathbf{h}}_{u_{n}^{\star}}\|^{2}$ [40]. This scheme represents a best-case reference in terms of beamforming capability under full CSIT. In contrast, the proposed ST-SIM architecture operates with only $N$ RF chains while exploiting a larger metasurface aperture with $V$ meta-atoms for wave-domain beamforming. For the proposed ST-SIM architecture, the overhead factor in (38) is
$\xi_{\mathrm{ST\text{-}SIM}}=1-\frac{M(N+1)}{L_{c}}$ (62)
whereas for the conventional MIMO benchmark we use
$\xi_{\mathrm{MIMO}}=1-\frac{2V}{L_{c}}$ (63)
where $L_{c}$ denotes the number of symbol periods in each channel coherence interval.
Figure 8: Objective function (44) of a DAL-aided and conventional SIM employing PC and AC layers versus $L_{\text{pc}}$ for $V\in\{9,16,25\}$ meta-atoms. Parameters are set to $Z=9$, $Q=25$, and $L_{\text{ac}}=4$. The case $Q=V=25$ corresponds to a baseline SIM architecture without absorbing meta-atoms. All metasurface layers are square. The signal from the RF chain first passes through the AC layers and, then, through the PC layers.
Figure 9 shows the time-averaged effective sum-rate versus the number of users $U$. We consider $V\in\{9,16\}$ and set $M$ according to (40). The PGD algorithm uses $\kappa_{\max}=10^{3}$, which definitely ensures that $\|{\mathbf{G}}_{0}-{\mathbf{G}}_{\mathrm{targ}}\|^{2}=10^{-8}$ for $V=9$ and $10^{-5}$ for $V=16$. As $U$ increases, the effective sum-rate of the proposed scheme grows for all values of $M$ and $V$. For $V=9$, the randomized ST-SIM outperforms the full-CSIT MIMO scheme for $U>100$, while for $V=16$ the crossover occurs around $U>500$. The gain arises from multiuser diversity enabled by channel randomization, whereas the digital benchmark benefits from perfect amplitude and phase knowledge at the transmitter at the cost of significantly higher feedback (see Remark 3).
Figure 9: TA effective sum-rate as a function of the number of users $U$ for three values of the number of time-slots $M$, with both $V=9$ and $V=16$ meta-atoms. Other ST-SIM parameters are $Q=576$, $Z=9$, $N=4$, $L_{\text{ac}}=2$, and $L_{\text{pc}}=6$. All metasurface layers are square. The signal from the RF chain first passes through the AC layers and, then, through the PC layers.
In order to assess the fairness gain of the randomized ST SIM-based beamforming relative to the conventional MIMO scheme, we adopt the following fairness index [45], averaged over $M$ time slots
which is presented in Fig. 10 as a function of the number of users $U$ for three different values of $M$. Results show that in the conventional MIMO scheme only $N=4$ users are scheduled on average for transmission, whereas the proposed randomized beamforming increases by a factor $M$ the average number of scheduled users thanks to its randomization-induced multiuser diversity. Larger values of $M$ enhance multiuser diversity but also increase training and CSI-feedback overheads (see Remark 2).
VII Conclusions
This paper introduced a beamforming framework for massive multiuser downlink connectivity based on randomized ST-SIM. By combining a time-varying DAL with multiple S-only metasurface layers, the architecture enables joint spatial–temporal wavefront shaping, increasing scheduling flexibility and exploiting multiuser diversity even under slowly varying channel conditions. A signal and propagation model was developed together with a PGD synthesis algorithm for configuring the metasurface coefficients. To reduce channel acquisition and feedback overhead, a partial-CSIT beamforming strategy based on randomized steering vectors and low-rate user feedback was proposed. Numerical results demonstrated that the randomized ST-SIM achieves competitive sum-rate performance with significantly reduced signaling requirements compared to full-CSIT solutions, supporting scalable downlink operation in dense networks.